A uniform ladder of weight W abd length 2a rests in limiting equilibrium with one end on a rough horizontal ground and the other end on a rough vertical wall. The coefficients of friction between the ladder and the ground and between the ladder and the wall are u and t respectively.
(a) Show that 5u + 6tu - 6 = 0
(b) Find the values of u and t, given that ut = 1/2
(c) Show by integration that tge centroid of a uniform semi-circular lamina of radius a from the centre is 4a/(3*pie).
1
Expert's answer
2021-02-05T11:48:28-0500

Explanations & Calculations


a)

  • For this consider the figure attached


  • Considering the vertical equilibrium,

"\qquad\qquad
\begin{aligned}
\small R+tS&-W=\small0\\
\small R+tS &= \small W\cdots(1)
\end{aligned}"

  • Fromt he horizi=ontal equilibrium,

"\qquad\qquad
\begin{aligned}
\small S-uR&= \small 0\\
\small S&= \small uR\cdots(2)
\end{aligned}"

  • By (1) and (2),

"\qquad\qquad
\begin{aligned}
\small R+t(uR)&= \small W\\
\small R &=\small \frac{W}{(1+ut)}
\end{aligned}"

  • To get a relationship in term only of friction coefficients, another equation is needed. Next is to consider of some torque.
  • Consider rightward torque about the point: ladder touches the wall,

"\qquad\qquad
\begin{aligned}
\small R(2a\cos\theta)-uR(2a\sin\theta)-W(a\cos\theta)&= \small 0\\
\small 2R(1-u\tan\theta)&= \small W\\
\small R&= \small \frac{W}{2(1-u\tan\theta)}
\end{aligned}"

  • Now removing R we get,

"\qquad\qquad
\begin{aligned}
\small \frac{1}{2(1-u\tan\theta)}&= \small \frac{1}{(1+ut)}\\
\small \tan\theta&= \small \frac{1-ut}{2u}
\end{aligned}"

  • Now, to get a relationship between u & t, "\small \tan\theta" should be known as it depends on the system.
  • Therefore, some data is missing from what you have provided.
  • Anyway when the "\small \theta=\tan^{-1}(\frac{5}{12})" , we get the needed result

"\qquad\qquad
\begin{aligned}
\small\frac{5}{12}&= \small \frac{1-ut}{2u}\\
\small 5u+6ut-6&= \small 0
\end{aligned}"

b)

  • When ut=1/2,

"\qquad\qquad
\begin{aligned}
\small5u+6(\frac{1}{2})-6&= \small0\\
\small u&= \small \frac{3}{5}(<1:\checkmark)\\
\\
\small t (\frac{3}{5})&= \small\frac{1}{2}\\
\small t&= \small \frac{5}{6}(<1:\checkmark)
\end{aligned}"

c)

  • For this consider the figure attached.


  • Consider an infinitesimal strip located at distance "\small x" from the center.
  • If the density of the lamina is "\rho", then we can write the follows.

"\qquad\qquad
\begin{aligned}
\small \delta m&= \small \delta A\times \rho\\
\small\delta m&= \small 2r\sin\theta\times\delta x\times\rho\\
\end{aligned}"

  • Mass of the lamina

"\qquad\qquad
\begin{aligned}
\small M&= \small \frac{\pi r^2}{2}\times \rho
\end{aligned}"

  • Since the lamina is homogeneous, the centroid should lye totally on the x axis. And if it lies at some "\small x'" distance from the center, we can write as follows for the system & each infinitesimal strips regarding the torque

"\qquad\qquad
\begin{aligned}
\small Mg\times x'&= \small \int\delta mg\times x\\
\small \frac{\pi r^2\rho}{2}g\times x'&= \small\int 2r\sin \theta.\rho .g. x.dx\cdots (1)\\


\end{aligned}"

  • To remove x & dx, we can write,

"\qquad\qquad
\begin{aligned}
\small x&= \small r\cos\theta\\
\small dx&= \small -r\sin\theta d\theta
\end{aligned}"

  • By (1),

"\qquad\qquad
\begin{aligned}
\small 
\end{aligned}" "\qquad\qquad
\begin{aligned}
\small \frac{\pi r^2\rho g}{2}x'&= \small -\int 2r^3\sin^2\theta\cos\theta .\rho.g.d\theta\\
\small x'&= \small \frac{-4r}{\pi}\int\sin^2\theta\cos\theta.d\theta\\

\end{aligned}"

  • Evaluating over "\large\frac{\pi}{2}"— 0,

"\qquad\qquad
\begin{aligned}
\small x'&= \small \frac{-4r}{\pi}\int_\frac{\pi}{2}^0\sin^2\theta .d(\sin\theta)\\
&= \small \frac{-4r}{\pi}\bigg[\frac{\sin^3\theta}{3}\big|_{\frac{\pi}{2}}^0\\
\small x'&= \small \frac{4r }{3\pi}
\end{aligned}"